# General Geospatial Inference with a Population Dynamics Foundation Model

### v5

@@ 1 Introduction
Models of population dynamics can be a powerful tool in understanding the impact of environmental, social, and economic factors on our well-being. Greater population level data could contribute to better planning and resource allocation, leading to improved outcomes for some of public health’s most vexing problems including noncommunicable disease, behavioral health disorders and climate related health impacts. There is evidence that having a deeper understanding of your postal code is a better predictor of long term health than your genetics code (Graham (2016)).
Whether we’re modeling the impact of an opioid epidemic or the population migration after an environmental disaster, such problems have traditionally needed dedicated teams and datasets to understand and model a single target variable (Gupta et al. (2003); Monnat et al. (2019)). To scale the capabilities and accessibility of these types of geospatial models, we introduce a novel Population Dynamics Foundation Model (PDFM) that uses machine learning to synthesize rich, globally available geospatial data such as maps, busyness, and aggregated search trends, along with environmental signals such as weather and air quality. To simplify the modeling of a broad range of problems affecting populations around the world, our approach allows these datasets to be at variable resolutions, consistency, and quality. For example, aggregated search trends might be insightful in heavily populated regions, while maps could be more useful in areas with lower internet penetration. We also show that incorporating embeddings from PDFM can imbue existing models with geospatial reasoning capabilities.
@@
  • SoTA Performance on Health, Socioeconomic, and Environment Geospatial Tasks. We benchmark the PDFM on interpolation, extrapolation, and super-resolution problems across 27 tasks (Sun et al. (2024)) covering health, socioeconomics, and the environment. Our foundation model provides SoTA performance in all 27 tasks for interpolation and 25 tasks for both extrapolation and super-resolution.
@@
  • Downstream Application to Forecasting Model. We show how the PDFM augments the SoTA forecasting foundation model TimesFM (Das et al. (2023)), to improve forecasts like unemployment at the county level and poverty at the postal code level (Figure 2). To accomplish this, we train a simple error correction model using PDFM embeddings and the TimesFM forecast for a single timestep. The resulting model surpasses a fully supervised forecasting method without fine-tuning any of the foundation models. Similar methods can be used to augment other existing geospatial classification and regression models with the PDFM embeddings.
  • Surpassing of Satellite Imagery and Geotagged Image Based Embeddings.Our experiments show that the PDFM embeddings surpass both the SatCLIP by Klemmer et al. (2023) and GeoCLIP by Vivanco Cepeda et al. (2024) embeddings in every downstream task except for tree cover extrapolation and elevation super-resolution. Previous results show a tradeoff where GeoCLIP outperformed satellite based approaches for human centric applications while PDFM is able to distill both human centric and built/natural environment signals.
@@ 2 Previous Work
Using machine learning approaches that leverage a variety of data sources has been shown to expand the spatiotemporal resolution and coverage of populations represented in geospatial maps. Deville et al. (2014) used a linear regression model with population-weighted least squares to study population movement over time using mobile phone data in Portugal and France. Web search trends data has been used to model automobile sales, unemployment claims, travel destination planning, and consumer confidence in Choi and Varian (2012) for countries around the world using linear models. Finely crafted web search features have been used in a regression model to predict influenza at the state level in Ginsberg et al. (2009). Rolf et al. (2021) used a featurization method over satellite imagery followed by task-specific regression to model forest cover, elevation, population density, nighttime lights, road length, and housing prices. A long short-term memory (LSTM) network was trained on precipitation, temperature, solar radiation, thermal radiation, snowfall, surface pressure, and basin attributes to forecast floods in Nearing et al. (2024). Lam et al. (2023) used a graph neural network to forecast weather using historical data globally. Each of these techniques provide unique abilities to model downstream geospatial tasks with strong performance that is actionable for policy makers and governments. However, each also required hand crafted features derived from each data source (mobile phone usage, web search, or satellite imagery) and in most cases, a custom model for each use case. This limits the ability to scale these types of models to new use cases and also increases the complexity of utilizing complementary datasets (for example, if we wanted to use both web search and mobile phone usage data).
@@
There have also been a variety of previous techniques that have sought to build a general geographic location encoding or geographic foundation model. Most focus on specific domains such as internet data, remote sensing imagery, or map databases. A semantic embedding for geo-coordinates was used to build GPS2Vec by Yin et al. (2019) using geotagged internet data including check-ins, tweets, and images. Satellite Contrastive Location-Image Pretraining (SatCLIP) Klemmer et al. (2023) is a “global, general-purpose geographic location encoder” based on two years of satellite imagery. SatCLIP outperforms a previous version called MOSAIKS (Rolf et al. (2021)) and a technique using contrastive self-supervision, CSP by Mai et al. (2023). GeoCLIP by Vivanco Cepeda et al. (2024) is an image-to-GPS model using a CLIP-inspired alignment between GPS and images from 4.72 million Flickr images. GeoCLIP often outperforms SatCLIP. We benchmarked our foundation model against both GeoCLIP and SatCLIP. A fine-tuned version of an LLM using OpenStreetMap by Manvi et al. (2023) was used to predict things such as population density, income, and home value. They focused on how performance varied with different language models and used a night lights dataset with GBDTs as a baseline. Our approach builds on this work by providing a wider range of general purpose datasets complementing human behavior signals (like those in GeoCLIP) with built and natural environment signals (like those in SatCLIP).
@@ 3 Experimental Results
We compared our model with inverse distance weighting (IDW) from Shepard (1968) because our experiments showed this to be the most competitive method for interpolation, extrapolation, and super-resolution across techniques that only made use of geographic coordinates. We also compared performance with SatCLIP and GeoCLIP because recent work shows these location encoders being state-of-the-art on environmental and socioeconomic tasks. We include tables breaking down performance by each of the individual datasets used in the PDFM to showcase the complementary information being modeled.
@@ 3.5 Forecasting
Our primary objective was to assess whether these embeddings could improve forecasts made for future time periods (six months for unemployment data and two years for poverty data). The results in Table 5 demonstrate a comparison between models that utilize PDFM embeddings and those that do not. The results in Table 5 show that for the mean absolute percentage error (MAPE) metric, the PDFM is able to improve on the TimesFM baseline performance and surpass the performance of ARIMA, suggesting that incorporating the PDFM can provide value without having to retrain existing forecasting models. We used a paired sample t-test to compare predictions from TimesFM + PDFM Embeddings and ARIMA, showing there to be a statistically significant difference (p-value < 0.025) when accounting for multiple comparison correction using Dunn (1961) in both poverty and unemployment.
@@ 5 Discussion
Our approach has some limitations. First, the temporal granularity, particularly for graph inputs, is nonideal and not aligned between covariates. For example, aggregated search trends were obtained over July of 2022 while maps data were obtained from May of 2024. Our results are promising despite this temporal mismatch, though a better understanding of temporal alignment and its effects could boost performance even further. Similarly, the current PDFM approach produces static embeddings; adding a time dimension to the graph where message passing is done over time could be an interesting future direction when data are available. The selection of input data used in the PDFM is not exhaustive and each dataset provides unique value. Future research could explore other diverse datasets, including those derived from public social media, census data, location-specific webpages, sales statistics, traffic flows, and news articles. Another limitation is the quality and source of our ground truth geospatial data. For example, the health target variables at the postal code and county level we used from CDC Places (Centers for Disease Control and Prevention (2024)) were derived from a small area estimation model using existing census data. Ideally, there would be higher quality and higher resolution (both temporally and spatially) data to better evaluate the performance on these categories. The PDFM incorporates both spatial proximity and feature similarity based graph edges. In future work we intend to explore more types of non-spatial edges. This would allow the nodes to learn through long distances that are not geographically connected but present similar patterns on some modalities. We also intend to explore new features that are available only to graph methods, such as edge only features. For example, going from Oakland to San Francisco involves a toll bridge only in one direction. Graphs can encode this asymmetric relationship that we have not fully explored with the labels we used. Further, datasets like aggregated mobility, capturing how people move across postal codes and counties, would bring interesting insights in epidemiological and health predictions via GNNs.
@@
Our work focuses on aggregating our input signals at relatively large spatial scales (postal code and county), longer temporal periods (on the order of months), and choosing the most common aggregated search trends or maps categories without retaining the actual queries or categories in our features. We have shown this feature design is generally applicable but it also helps to preserve privacy. In future scenarios where the PDFM framework incorporates more sensitive signals mapping sparser populations, techniques such as differential privacy Bavadekar et al. (2021); Kaplan et al. (2019) will need to be applied to provide clear privacy guarantees.
@@ 6.1 Datasets
Aggregated Search Trends. We computed the aggregate counts of the top 500 queries that were searched at least 20 times during July of 2022 in each postal code (Sun et al. (2024)). This resulted in over 1MM unique queries. We ranked the queries by their countrywide popularity, represented by the total number of postal codes that each query appeared in. We selected the top 1,000 queries from this list as a representation of aggregated search trends activity at the postal code level across the country. The lowest ranked or least common query appeared in over 700 postal codes. In each postal code we built our 1,000 dimensional feature vector with its respective frequencies for each query and scaled the vector to sum to 100. We designed these features to preserve privacy by ensuring we selected the most common queries. Each of the 1,000 queries occurred at least 14,000 times in the dataset and the query texts are ablated from subsequent analysis. The resulting dataset contained no personally identifiable information.
@@
Remote Sensing. We incorporated satellite imagery embeddings derived from the ViT16-L40 version of SatCLIP (Klemmer et al. (2023)) using the centroid of each postal code as the index to each embedding. The SatCLIP model aims to be a global, general-purpose geographic location encoder and summarizes 100,000 tiles of Sentinel-2 satellite imagery captured between January 1, 2021 and May 17, 2023.
@@ 6.2.1 Graph Neural Network Foundation Model (Phase 1)
Subgraph Sampling. We perform subgraph sampling to create subgraphs for training large-scale GNNs and adding stochasticity to the model as recommended by the GraphSAGE framework introduced in Hamilton et al. (2017). For each geographical node (treated as the seed node), we sample both postal code and county nodes up to four hops away from the seed node. Specifically, we start from the seed node, traverse in a breadth-first fashion for each edge set, sample a fixed number of nodes in a weighted manner, and terminate when we’re four hops away from the seed node. This approach leads to a total number of subgraphs that is equal to the total number of postal code and county nodes. Each of these subgraphs represents the local neighborhood associated with its corresponding seed node. For modeling, each subgraph and the corresponding label serve as a single data point for training and validation purposes.
@@ 6.2.2 Downstream Models (Phase 2)
Linear Regression. For each task and label combination, we trained a Ridge regression model with default configurations from the popular Scikit-learn library (Fabian (2011)).
Multilayer Perceptron (MLP). MLP is a type of feedforward neural network suitable for modeling nonlinear relationships between variables. For each task and label combination, we used the TensorFlow (Abadi et al. (2016)) library to train a three layer MLP with rectified linear (ReLU) activation and a 20% dropout rate, using the Adam optimizer with an initial learning rate of 0.005 and cosine decay over 40 epochs. The layer dimensions are 512, 256, and 128.
Gradient Boosted Decision Trees (GBDT). GBDT is a common state of the art method for modeling tabular data (Shwartz-Ziv and Armon (2022); Grinsztajn et al. (2022)). We used the GBDT implementation from the popular LightGBM (Ke, 2017) library with lightly tuned hyperparameters: up to 3000 trees with a maximum of 31 leaves per tree and a minimum of 40 samples per leaf, trained with a learning rate of 0.02. Each model trained in a few minutes on a single CPU and typically performed better than linear regression.
@@ 6.3.1 Problem Definition
Missing data is a common issue with surveys such as the census that might miss certain demographics, locations, or answers (Brick and Kalton (1996)). The goal of imputation via spatial interpolation is to generate the most likely data to substitute those missing values.
@@ 6.3.2 Benchmark Datasets
We used an existing general geospatial benchmark by (Sun et al. (2024)), retrieved from Data Commons (2024) and the Earth Engine catalog (Gorelick et al. (2017)). Our health tasks are all of the 2022 health related metrics from CDC Places that were available in Data Commons at the postal code resolution. We selected all socioeconomic and environmental tasks that were used by Rolf et al. (2021) and available in Data Commons or the Earth Engine Catalog. These included income, home value, night lights, tree cover, and elevation. Klemmer et al. (2023) did a comparison of median income, home value, elevation, and population across a range of geolocation encoders. They found SatCLIP and GeoCLIP to be the optimal encodings. We evaluated both SatCLIP and GeoCLIP across all of our variables in our experiments.
Tkachenko et al. (2017) studied diabetes by correlating specific queries with candidate risk variables. Cesare et al. (2019) used search trends related to physical activity and diet along with social media, demographics, and built environment variables to predict 2013 county level obesity prevalence across the US using a linear model with $R^{2}$ of 0.624 and 0.55 for females and males respectively. Smoking, binge drinking, stroke, coronary heart disease, obesity, diabetes, and poor physical health in 208 Designated Market Areas (DMAs) across the US by selecting categories of search trends related to these tasks (Jaidka et al. (2021)). We build on these examples showing how we are able to improve performance by incorporating complementary datasets for these types of health tasks.

### v6

@@ 1 Introduction
Models of population dynamics can be a powerful tool in understanding the impact of environmental, social, and economic factors on our well-being. Greater population level data could contribute to better planning and resource allocation, leading to improved outcomes for some of public health’s most vexing problems including noncommunicable disease, behavioral health disorders and climate related health impacts. There is evidence that having a deeper understanding of your postal code is a better predictor of long term health than your genetics code (graham2016your).
Whether we’re modeling the impact of an opioid epidemic or the population migration after an environmental disaster, such problems have traditionally needed dedicated teams and datasets to understand and model a single target variable (gupta2003uses, monnat2019using). To scale the capabilities and accessibility of these types of geospatial models, we introduce a novel Population Dynamics Foundation Model (PDFM) that uses machine learning to synthesize rich, globally available geospatial data such as maps, busyness, and aggregated search trends, along with environmental signals such as weather and air quality. To simplify the modeling of a broad range of problems affecting populations around the world, our approach allows these datasets to be at variable resolutions, consistency, and quality. For example, aggregated search trends might be insightful in heavily populated regions, while maps could be more useful in areas with lower internet penetration. We also show that incorporating embeddings from PDFM can imbue existing models with geospatial reasoning capabilities.
@@
  • SoTA Performance on Health, Socioeconomic, and Environment Geospatial Tasks. We benchmark the PDFM on interpolation, extrapolation, and super-resolution problems across 27 tasks (sun2024community) covering health, socioeconomics, and the environment. Our foundation model provides SoTA performance in all 27 tasks for interpolation and 25 tasks for both extrapolation and super-resolution.
@@
  • Downstream Application to Forecasting Model. We show how the PDFM augments the SoTA forecasting foundation model TimesFM (das2023decoder), to improve forecasts like unemployment at the county level and poverty at the postal code level (Figure 2). To accomplish this, we train a simple error correction model using PDFM embeddings and the TimesFM forecast for a single timestep. The resulting model surpasses a fully supervised forecasting method without fine-tuning any of the foundation models. Similar methods can be used to augment other existing geospatial classification and regression models with the PDFM embeddings.
  • Surpassing of Satellite Imagery and Geotagged Image Based Embeddings.Our experiments show that the PDFM embeddings surpass both the SatCLIP by klemmer2023satclip and GeoCLIP by cepeda2024geoclip embeddings in every downstream task except for tree cover extrapolation and elevation super-resolution. Previous results show a tradeoff where GeoCLIP outperformed satellite based approaches for human centric applications while PDFM is able to distill both human centric and built/natural environment signals.
@@ 2 Previous Work
Using machine learning approaches that leverage a variety of data sources has been shown to expand the spatiotemporal resolution and coverage of populations represented in geospatial maps. deville2014dynamic used a linear regression model with population-weighted least squares to study population movement over time using mobile phone data in Portugal and France. Web search trends data has been used to model automobile sales, unemployment claims, travel destination planning, and consumer confidence in choi2012predicting for countries around the world using linear models. Finely crafted web search features have been used in a regression model to predict influenza at the state level in ginsberg2009detecting. rolf2021generalizable used a featurization method over satellite imagery followed by task-specific regression to model forest cover, elevation, population density, nighttime lights, road length, and housing prices. A long short-term memory (LSTM) network was trained on precipitation, temperature, solar radiation, thermal radiation, snowfall, surface pressure, and basin attributes to forecast floods in nearing2024global. lam2023learning used a graph neural network to forecast weather using historical data globally. Each of these techniques provide unique abilities to model downstream geospatial tasks with strong performance that is actionable for policy makers and governments. However, each also required hand crafted features derived from each data source (mobile phone usage, web search, or satellite imagery) and in most cases, a custom model for each use case. This limits the ability to scale these types of models to new use cases and also increases the complexity of utilizing complementary datasets (for example, if we wanted to use both web search and mobile phone usage data).
@@
There have also been a variety of previous techniques that have sought to build a general geographic location encoding or geographic foundation model. Most focus on specific domains such as internet data, remote sensing imagery, or map databases. A semantic embedding for geo-coordinates was used to build GPS2Vec by yin2019gps2vec using geotagged internet data including check-ins, tweets, and images. Satellite Contrastive Location-Image Pretraining (SatCLIP) klemmer2023satclip is a “global, general-purpose geographic location encoder” based on two years of satellite imagery. SatCLIP outperforms a previous version called MOSAIKS (rolf2021generalizable) and a technique using contrastive self-supervision, CSP by mai2023csp. GeoCLIP by cepeda2024geoclip is an image-to-GPS model using a CLIP-inspired alignment between GPS and images from 4.72 million Flickr images. GeoCLIP often outperforms SatCLIP. We benchmarked our foundation model against both GeoCLIP and SatCLIP. A fine-tuned version of an LLM using OpenStreetMap by manvi2023geollm was used to predict things such as population density, income, and home value. They focused on how performance varied with different language models and used a night lights dataset with GBDTs as a baseline. Our approach builds on this work by providing a wider range of general purpose datasets complementing human behavior signals (like those in GeoCLIP) with built and natural environment signals (like those in SatCLIP).
@@ 3 Experimental Results
We compared our model with inverse distance weighting (IDW) from shepard1968two because our experiments showed this to be the most competitive method for interpolation, extrapolation, and super-resolution across techniques that only made use of geographic coordinates. We also compared performance with SatCLIP and GeoCLIP because recent work shows these location encoders being state-of-the-art on environmental and socioeconomic tasks. We include tables breaking down performance by each of the individual datasets used in the PDFM to showcase the complementary information being modeled.
@@ 3.5 Forecasting
Our primary objective was to assess whether these embeddings could improve forecasts made for future time periods (six months for unemployment data and two years for poverty data). The results in Table 5 demonstrate a comparison between models that utilize PDFM embeddings and those that do not. The results in Table 5 show that for the mean absolute percentage error (MAPE) metric, the PDFM is able to improve on the TimesFM baseline performance and surpass the performance of ARIMA, suggesting that incorporating the PDFM can provide value without having to retrain existing forecasting models. We used a paired sample t-test to compare predictions from TimesFM + PDFM Embeddings and ARIMA, showing there to be a statistically significant difference (p-value < 0.025) when accounting for multiple comparison correction using dunn1961multiple in both poverty and unemployment.
@@ 5 Discussion
Our approach has some limitations. First, the temporal granularity, particularly for graph inputs, is nonideal and not aligned between covariates. For example, aggregated search trends were obtained over July of 2022 while maps data were obtained from May of 2024. Our results are promising despite this temporal mismatch, though a better understanding of temporal alignment and its effects could boost performance even further. Similarly, the current PDFM approach produces static embeddings; adding a time dimension to the graph where message passing is done over time could be an interesting future direction when data are available. The selection of input data used in the PDFM is not exhaustive and each dataset provides unique value. Future research could explore other diverse datasets, including those derived from public social media, census data, location-specific webpages, sales statistics, traffic flows, and news articles. Another limitation is the quality and source of our ground truth geospatial data. For example, the health target variables at the postal code and county level we used from CDC Places (cdcplaces2024) were derived from a small area estimation model using existing census data. Ideally, there would be higher quality and higher resolution (both temporally and spatially) data to better evaluate the performance on these categories. The PDFM incorporates both spatial proximity and feature similarity based graph edges. In future work we intend to explore more types of non-spatial edges. This would allow the nodes to learn through long distances that are not geographically connected but present similar patterns on some modalities. We also intend to explore new features that are available only to graph methods, such as edge only features. For example, going from Oakland to San Francisco involves a toll bridge only in one direction. Graphs can encode this asymmetric relationship that we have not fully explored with the labels we used. Further, datasets like aggregated mobility, capturing how people move across postal codes and counties, would bring interesting insights in epidemiological and health predictions via GNNs.
@@
Our work focuses on aggregating our input signals at relatively large spatial scales (postal code and county), longer temporal periods (on the order of months), and choosing the most common aggregated search trends or maps categories without retaining the actual queries or categories in our features. We have shown this feature design is generally applicable but it also helps to preserve privacy. In future scenarios where the PDFM framework incorporates more sensitive signals mapping sparser populations, techniques such as differential privacy bavadekar2021google, kaplan2019differentially will need to be applied to provide clear privacy guarantees.
@@ 6.1 Datasets
Aggregated Search Trends. We computed the aggregate counts of the top 500 queries that were searched at least 20 times during July of 2022 in each postal code (sun2024community). This resulted in over 1MM unique queries. We ranked the queries by their countrywide popularity, represented by the total number of postal codes that each query appeared in. We selected the top 1,000 queries from this list as a representation of aggregated search trends activity at the postal code level across the country. The lowest ranked or least common query appeared in over 700 postal codes. In each postal code we built our 1,000 dimensional feature vector with its respective frequencies for each query and scaled the vector to sum to 100. We designed these features to preserve privacy by ensuring we selected the most common queries. Each of the 1,000 queries occurred at least 14,000 times in the dataset and the query texts are ablated from subsequent analysis. The resulting dataset contained no personally identifiable information.
@@
Remote Sensing. We incorporated satellite imagery embeddings derived from the ViT16-L40 version of SatCLIP (klemmer2023satclip) using the centroid of each postal code as the index to each embedding. The SatCLIP model aims to be a global, general-purpose geographic location encoder and summarizes 100,000 tiles of Sentinel-2 satellite imagery captured between January 1, 2021 and May 17, 2023.
@@ 6.2.1 Graph Neural Network Foundation Model (Phase 1)
Subgraph Sampling. We perform subgraph sampling to create subgraphs for training large-scale GNNs and adding stochasticity to the model as recommended by the GraphSAGE framework introduced in hamilton2017inductive. For each geographical node (treated as the seed node), we sample both postal code and county nodes up to four hops away from the seed node. Specifically, we start from the seed node, traverse in a breadth-first fashion for each edge set, sample a fixed number of nodes in a weighted manner, and terminate when we’re four hops away from the seed node. This approach leads to a total number of subgraphs that is equal to the total number of postal code and county nodes. Each of these subgraphs represents the local neighborhood associated with its corresponding seed node. For modeling, each subgraph and the corresponding label serve as a single data point for training and validation purposes.
@@ 6.2.2 Downstream Models (Phase 2)
Linear Regression. For each task and label combination, we trained a Ridge regression model with default configurations from the popular Scikit-learn library (fabian2011scikit).
Multilayer Perceptron (MLP). MLP is a type of feedforward neural network suitable for modeling nonlinear relationships between variables. For each task and label combination, we used the TensorFlow (abadi2016tensorflow) library to train a three layer MLP with rectified linear (ReLU) activation and a 20% dropout rate, using the Adam optimizer with an initial learning rate of 0.005 and cosine decay over 40 epochs. The layer dimensions are 512, 256, and 128.
Gradient Boosted Decision Trees (GBDT). GBDT is a common state of the art method for modeling tabular data (shwartz2022tabular, grinsztajn2022why). We used the GBDT implementation from the popular LightGBM (Ke, 2017) library with lightly tuned hyperparameters: up to 3000 trees with a maximum of 31 leaves per tree and a minimum of 40 samples per leaf, trained with a learning rate of 0.02. Each model trained in a few minutes on a single CPU and typically performed better than linear regression.
@@ 6.3.1 Problem Definition
Missing data is a common issue with surveys such as the census that might miss certain demographics, locations, or answers (brick1996handling). The goal of imputation via spatial interpolation is to generate the most likely data to substitute those missing values.
@@ 6.3.2 Benchmark Datasets
We used an existing general geospatial benchmark by (sun2024community), retrieved from datacommons2024 and the Earth Engine catalog (gorelick2017google). Our health tasks are all of the 2022 health related metrics from CDC Places that were available in Data Commons at the postal code resolution. We selected all socioeconomic and environmental tasks that were used by rolf2021generalizable and available in Data Commons or the Earth Engine Catalog. These included income, home value, night lights, tree cover, and elevation. klemmer2023satclip did a comparison of median income, home value, elevation, and population across a range of geolocation encoders. They found SatCLIP and GeoCLIP to be the optimal encodings. We evaluated both SatCLIP and GeoCLIP across all of our variables in our experiments.
tkachenko2017google studied diabetes by correlating specific queries with candidate risk variables. cesare2019use used search trends related to physical activity and diet along with social media, demographics, and built environment variables to predict 2013 county level obesity prevalence across the US using a linear model with $R^{2}$ of 0.624 and 0.55 for females and males respectively. Smoking, binge drinking, stroke, coronary heart disease, obesity, diabetes, and poor physical health in 208 Designated Market Areas (DMAs) across the US by selecting categories of search trends related to these tasks (jaidka2021information). We build on these examples showing how we are able to improve performance by incorporating complementary datasets for these types of health tasks.